# Empirical Orlicz norms

### v2

@@ 3 Central limit theorem
Example (Exponential distribution)
@@
Example (Weibull distribution)
@@
Example (Normal distribution)
@@
This highlights that the empirical Orlicz norm does not admit any parametric rate of convergence uniformly on the class $\{X:\|X\|_{\psi}\leq 1\}$ of all distributions with bounded Orlicz norm. Note that Theorem 3.5 does not rule out a potential uniform logarithmic rate of convergence. However, we conjecture that there does not exist any uniform rate of convergence for the empirical Orlicz norm.
@@
Proof: No rate of convergence (Theorem 3.5)
@@
However, in contrast to Theorem 3.1, the addends $\psi_{\alpha}(X_{i})$ comprising $G_{n}(1)$ have regularly varying upper tails with exponent $\gamma$, and finite lower tails. Thus, the generalized central limit theorem (Feller,, 1971, IX.8) yields convergence of $n^{1-\frac{1}{\gamma}}[G_{n}(1)-\mathbb{E}G_{n}(1)]$ towards a fully right-skewed $\gamma$-stable random variable. In particular, the estimation error $\widehat{\sigma}_{\psi_{\alpha}}-\sigma_{\psi_{\alpha}}$ is of exact order $n^{\frac{1}{\gamma}-1}$. Since $\gamma\in(1,2)$ is arbitrary, any rate $n^{\beta}$ with $\beta\in(0,\frac{1}{2})$ can be obtained.

### v3

@@ 3 Central limit theorem
Example (Exponential distribution).
@@
Example (Weibull distribution).
@@
Example (Normal distribution).
@@
This highlights that the empirical Orlicz norm does not admit any parametric rate of convergence uniformly on the class $\mathcal{X}_{\psi}=\{X:\|X\|_{\psi}\leq 1\}$ of all distributions with bounded Orlicz norm. Note that Theorem 3.5 does not rule out a potential uniform logarithmic rate of convergence. However, we conjecture that there does not exist any uniform rate of convergence for the empirical Orlicz norm, as the following statistical lower bound holds for any estimator.
Theorem 3.6.
For any Orlicz function $\psi$ and any estimator $\widetilde{\sigma}_{\psi}=\widetilde{\sigma}_{\psi}(X_{1},\ldots,X_{n})$ with $X_{1},\ldots,X_{n}\overset{iid}{\sim}X$, and any rate $r_{n}\to 0$,
$\limsup_{n\to\infty}\sup_{X\in\mathcal{X}_{\psi}}P\left(\widetilde{\sigma}_{\psi}(X_{1},\ldots,X_{n})-\|X\|_{\psi}>r_{n}\right)=1.$
The claim of Theorem 3.6 is stronger in that it also excludes a slower-than-polynomial rate for any estimator, but requires a stricter form of uniformity than Theorem 3.5. Specifically, in Theorem 3.5, the distribution of $X$ is fixed for all $n$, while Theorem 3.6 allows for $X$ to vary with $n$. We also note that in any parametric family of distributions, the Orlicz norm can be estimated via plug-in of a parameter estimator, which typically converges at rate $\sqrt{n}$. The empirical Orlicz norm is, on the other hand, a model-free estimator.
Example.
Data-driven estimates of the probability $\overline{F}(t)=P(X>t)$, for very large $t$ are typically derived via extreme value theory, and applied, for example, in hydrology (Katz et al.,, 2002). This mathematical approach is able to yield accurate probability estimates for any $t\leq t_{n}$ in the regime $\overline{F}(t_{n})\gg\exp(-\sqrt{n})$, see de Haan and Ferreira, (2006, Thm. 4.4.1). As an alternative to the extreme value methodology, the empirical Orlicz norm can be plugged into the conservative tail inequality $\overline{F}(t)\leq\overline{F}^{*}(t)=1/\psi(t/\|X\|_{\psi})$ to obtain an empirical upper bound $\overline{F}^{\star}(t)=1/\psi(t/\widehat{\sigma}_{\psi})$, which upholds for very large values of $t$. Specifically, rainfall measurements are often described as having sub-Weibull tails (Papalexiou and Koutsoyiannis,, 2013; Marra et al.,, 2023), which naturally leads to the use of sub-Weibull norms $\sigma_{\psi_{\alpha}}=\|X\|_{\psi_{\alpha}}$. The tail bound takes the form
$P(X>t)$ $\leq\overline{F}^{*}(t)=\frac{1}{\exp(t^{\alpha}/\sigma_{\psi_{\alpha}}^{\alpha})-1}\leq\frac{1}{\exp(t^{\alpha}/\widehat{\sigma}_{\psi_{\alpha}}^{\alpha})-1}\Delta_{n}(t)=\overline{F}^{\star}(t)\Delta_{n}(t),$
where, for some $\overline{\sigma}$ between $\sigma_{\psi_{\alpha}}$ and $\widehat{\sigma}_{\psi_{\alpha}}$,
$\left|\Delta_{n}(t)-1\right|$ $\leq\left|\frac{\exp(t^{\alpha}/\widehat{\sigma}_{\psi_{\alpha}}^{\alpha})-\exp(t^{\alpha}/{\sigma}_{\psi_{\alpha}}^{\alpha})}{\exp(t^{\alpha}/\sigma_{\psi_{\alpha}}^{\alpha})-1}\right|\quad\leq\frac{\alpha t^{\alpha}\left|\widehat{\sigma}_{\psi_{\alpha}}-\sigma_{\psi_{\alpha}}\right|}{\overline{\sigma}^{\alpha+1}}\frac{\exp(t^{\alpha}/\overline{\sigma}^{\alpha})}{\exp(t^{\alpha}/\sigma_{\psi_{\alpha}}^{\alpha})-1}$
$\leq\frac{\alpha t^{\alpha}|\widehat{\sigma}_{\psi_{\alpha}}-\sigma_{\psi_{\alpha}}|}{\overline{\sigma}^{\alpha+1}}\frac{\exp(t^{\alpha}/\sigma_{\psi_{\alpha}}^{\alpha})}{\exp(t^{\alpha}/\sigma_{\psi_{\alpha}}^{\alpha})-1}\exp\left(\tfrac{\alpha t^{\alpha}}{\overline{\sigma}^{\alpha+1}}|\widehat{\sigma}_{\psi_{\alpha}}-\sigma_{\psi_{\alpha}}|\right)$
We thus find that $\Delta_{n}(t)\to 1$ uniformly for $t\in[0,t_{n}]$ if $t_{n}\ll|\widehat{\sigma}_{\psi_{\alpha}}-{\sigma}_{\psi_{\alpha}}|^{1/\alpha}$. Hence, the rate of convergence of $\widehat{\sigma}_{\psi}$ governs how far we can reliably extrapolate into the tail. Under the conditions of Theorem 3.1, the rate implies the bound $t_{n}\ll n^{2/\alpha}$, corresponding to a tail probability bound of approximately $\overline{F}^{\star}(t_{n})=\exp(-o(n^{2}))$. In comparison to the extreme value approach, we obtain an upper bound instead of an exact estimate, but the former is reliable way farther into the tail.
@@
Proof: No rate of convergence
Proof of Theorem 3.5.
@@
However, in contrast to Theorem 3.1, the addends $\psi_{\alpha}(X_{i})$ comprising $G_{n}(1)$ have regularly varying upper tails with exponent $\gamma$, and finite lower tails. Thus, the generalized central limit theorem (Feller,, 1971, IX.8) yields convergence of $n^{1-\frac{1}{\gamma}}[G_{n}(1)-\mathbb{E}G_{n}(1)]$ towards a fully right-skewed $\gamma$-stable random variable. In particular, the estimation error $\widehat{\sigma}_{\psi_{\alpha}}-\sigma_{\psi_{\alpha}}$ is of exact order $n^{\frac{1}{\gamma}-1}$. Since $\gamma\in(1,2)$ is arbitrary, any rate $n^{\beta}$ with $\beta\in(0,\frac{1}{2})$ can be obtained. ∎
Proof of Theorem 3.6.
Let $X_{1},\ldots,X_{n}\sim X=0$ almost surely, and $Y_{1}^{n},\ldots,Y_{n}^{n}\sim Y^{n}$ where $P(Y^{n}=\psi^{-1}(n^{2}))=1/n^{2}$ and $P(Y^{n}=0)=1-1/n^{2}$. Then $\|X\|_{\psi}=0$ and $\|Y^{n}\|_{\psi}=1$, but any estimator satisfies
$P\left(\widetilde{\sigma}_{\psi}(X_{1},\ldots,X_{n})=\widetilde{\sigma}_{\psi}(Y_{1}^{n},\ldots,Y_{n}^{n})\right)\;\geq\;P\left((X_{1},\ldots,X_{n})=(Y_{1}^{n},\ldots,Y_{n}^{n})\right)\geq 1-\tfrac{1}{n}.$
Hence, the estimator can not uniformly distinguish the cases $\sigma_{\psi}=0$ and $\sigma_{\psi}=1$. ∎